# Modelling Conduction Cooling of Superconducting Accelerator Magnets using a Thermal Thin Shell Approximation

### v1

@@ Abstract
Understanding the thermal behaviour of superconducting accelerator magnets is essential to ensure stable and reliable operation. This work presents an extension of the Finite Element Quench Simulator (FiQuS) Multipole module to include collar and pole regions of accelerator magnets, which influences the overall thermal response. A thermal thin shell approximation (TSA), which is shown to be effective in previous works, is employed to model thermal insulation layers efficiently, replacing an insulation surface mesh. The main novelty of this work lies in the development of a method to model the thermal connection between the magnet winding and the collar and pole regions via the TSA.
To assess the accuracy and computational efficiency of the novel method, temperature and field variations are computed for a current ramp-up scenario. The thermal solution is coupled to a fully resolved magnetodynamic solution to capture the interaction between thermal and electromagnetic behaviour. The results obtained with the TSA are then compared to classical finite element (FE) solutions with explicitly meshed insulation domains. The TSA predicts the maximum temperature within 2–4 % of the reference solution while substantially reducing mesh complexity and achieving up to a 5 times speed-up in computation time. While the TSA has traditionally been employed for short-duration quench simulations with high heat fluxes between magnet turns, these results demonstrate its reliability and efficiency for current ramp scenarios with low heat fluxes, significantly expanding its application range beyond what has been previously reported in the literature.
@@ 1 Introduction
The thermal stability of magnets is of critical importance for superconducting accelerator components. At the European Council for Nuclear Research (CERN), the Simulation of Transient Effects in Accelerator Magnets (STEAM) [1, 2] framework provides a collection of tools to model the transient behaviour of these magnets. Within this framework, the Finite Element Quench Simulator (FiQuS) [3] is a coupled magnetodynamic-thermal finite-element based tool for a variety of magnets types (e.g. high-temperature superconducting pancake coils [4], multipole magnets [3], canted cosine-theta magnets [5]). FiQuS relies on the Gmsh Python application programming interface [6] for geometry and mesh creation, and GetDP [7, 8, 9] for solving the governing equations.
Quench simulations typically involve large temperature gradients between adjacent half-turns, as the thermal conductivity of the superconducting cable is significantly larger than that of the surrounding electric insulation. Capturing these gradients accurately typically requires a high-quality, fully resolved mesh of the insulation layers [10, Sec. 5.3], leading to a large number of degrees of freedom (DoFs) and, consequently, high computational costs. For coupled FE simulations, two different meshes are often used for the thermal and magnetic models [11]. By enforcing suitable interface conditions, the thermal thin shell approximation (TSA) proposes an efficient alternative without the need for explicit meshing [11, 12, 13, 14, 15].
Figure 1: Illustration of the electromagnetic computational domain $\Omega_{\mathrm{EM}}$ and thermal computational domain $\Omega_{\mathrm{TH}}$ for a 12 T dipole magnet design suggested in [16]. Only a quarter of the magnet is shown. The collar, pole, and ground insulation regions have been newly integrated into FiQuS. The materials are specified in Tab. 1.
While commercial tools like COMSOL Multiphysics can model thin thermal layers [17], these implementations rely on proprietary, closed-source mechanisms. The detailed derivation of the TSA is presented and implemented in an open-source framework in [13]. Despite these advances, current state-of-the-art models still typically neglect the thermal influence of structural components such as the pole and collar.
Cooling mechanisms are critical for stable magnet operation. Standard approaches for modelling immersive cooling methods include convective cooling boundary conditions via a heat transfer coefficient [18], e.g. in FiQuS [19, Sec. 2.2], or by modelling a confined helium layer next to the insulation, as is done in ANSYS [20]. In the Large Hadron Collider (LHC) main dipoles, the non-impregnated Nb-Ti cables allow superfluid helium to permeate the coil and provide efficient cooling [21]. However, the amount of liquid helium required and potential leakages can lead to a high cost.
@@ 2.1 Magneto-thermal formulation
Magnetodynamic equations are considered in the electromagnetic computational domain $\Omega_{\mathrm{EM}}$ in an $\mathbf{A}$-formulation. The magnetic vector potential $\mathbf{A}$ is related to the magnetic flux density $\mathbf{B}~=~\nabla\times\mathbf{A}$ and the electric field $\mathbf{E}~=~-\tfrac{\partial~\mathbf{A}}{\partial~t}~+~\nabla~\phi$. The electric scalar potential $\phi$ can be eliminated by a gauge choice: $\mathbf{A}^{*}~=~\mathbf{A}~+~\int_{t_{0}}^{t}~\nabla\phi\,~\mathrm{dt}$, leading to the strong $\mathbf{A}^{*}$-formulation with the governing differential equation [26] [27, Sec. 5.18]:
@@
where $\kappa=\kappa(|\mathbf{B}|,~T)$ denotes the thermal conductivity, $C_{V}=C_{V}(T)$ the volumetric heat capacity, $\vec{n}$ the normal to the boundary, and $Q_{\mathrm{ht}}$ the additional Joule losses in the half-turns, which are triggered when the local current density exceeds the current sharing threshold. This corresponds to Eq. (3) in [12]. The power density $P=\sigma|\mathbf{E}|^{2}$ is a modified, ramp rate dependent heat source, introduced to include a non-zero heat source below current sharing regime. Both the heat transfer coefficient $h=h(T)$ and the temperature of the liquid helium $T_{\mathrm{He}}=1.9$ K are used in Eq. (3) to define a Robin-type boundary condition for the cooling holes. The last equation imposes adiabatic conditions, i.e., no heat can escape through $\partial\Omega_{\mathrm{TH}}\!\setminus\!\Gamma_{\mathrm{Cool}}$. The half-turns region is modelled as a single homogeneous domain, where the different material properties are accounted for via volumetric fractions $f$, e.g. for $\sigma|_{\Omega_{\mathrm{ht}}}=f_{\mathrm{{}\mathrm{Cu}}}\sigma_{{}\mathrm{Cu}}$, as proposed in [12, Eqs. (1), (2) and (4)].
@@ 2.2.1 General idea and closest neighbour mapping
Figure 4: Illustration of a free closest neighbour map and an enforced map. For clarity, only the collar temperature is shown at a specific point in time, for a quenching simulation and a stainless steel collar. Note the difference in local temperatures.
@@
For quench simulations or when using a low thermal conducting collar material, this closest neighbouring mesh node mapping might cause numerical artifacts in the simulation. This is illustrated by Fig. 4. On the left figure, two links are associated with the same mesh node, and since the total heat flux is conserved, there is an increase in local temperature compared to the figure on the right. This can lead to large discrepancies in maximum collar temperatures. The problem can be avoided by applying a mortaring method to link both sides of a non-conformal mesh, enforcing equal heat flux by adding Lagrange multipliers [33]. This comes at the cost of additional unknowns and consequently a higher computational complexity.
Instead, a small mesh manipulation is implemented which enforces the mesh node position on the collar side to guarantee a one-to-one mapping. As illustrated by Fig. 4, this removes the artificial local differences in temperature, compromising in the collar mesh freedom.
@@ 2.2.2 Appropriate length scaling
As mentioned in the previous section, the TSA assumes that the tangential heat flux between different TSLs in the insulation layer is negligible. For simulations with fast quenching half-turns, this assumption leads to accurately computed maximum temperatures [12]. However, one must be cautious when generalising this. First of all, the average insulation thickness between the collar and half-turns or wedges (ground insulation in Fig. 1) is considerably larger than the in-between half-turns and wedges (cable insulation in Fig. 1). As a result, more internal discretization levels in the TSA are needed. Moreover, the half-turn width ($w_{\mathrm{ht}}$ in Fig. 1) is typically much smaller than its height ($h_{\mathrm{ht}}$ in Fig. 4).
Taking into account the cable insulation with a thickness $d$, we make a distinction between lines along the radial and azimuthal direction. The former lines, appearing between one half-turn and the collar or pole, are assumed to have a contact length of about $15\%$ larger ($h_{\mathrm{ht}}+2d]/h_{\mathrm{ht}}=1.1508$, see Fig. 4) for a 12 T dipole magnet. Similarly, the azimuthal lines, only appearing at the poles, represent a contact length which is $7\%$ larger ($w_{\mathrm{pole}}~/~w_{\mathrm{ht}}~=~1.0701$, see Fig. 1).
We use a scaling factor in order to diminish the error introduced by these geometric differences for the thin shell length, without compromising the computational speed by adding additional calculations or restraining the mesh. A similar scaling of material properties has previously been applied in [15] for pancake coils. In our case, we model the TSL as an extension of the half-turns (thus having the bare half-turn length) and we scale the material parameters with a factor, effectively modelling a longer thin shell, as depicted in Fig. 4. The main benefit of keeping the thin shell length the same as the half-turns on a geometrical level is for the mapping and for the mesh enforcement explained earlier. Concerning the other TSLs connecting the wedges to the collar and pole regions, no correction is applied, assuming the insulation thickness is very thin compared to the wedge or half-turn length. This scaling law can only partially cover the effect of the tangential heat flow and it does not account for the mid-pole additional insulation. Moreover, since this linear scaling is applied to non-linear material properties, it cannot match the reference model perfectly. However, the approach leads to reasonable accuracy as shown in Sec. 3.
@@ 3.1.1 Fully resolved mesh models
The mesh quality is analysed by varying the mesh sizes, i.e. the characteristic size of the mesh triangles, and comparing the simulation results to those obtained with the most refined mesh. We consider variation of the insulation mesh and collar mesh sizes. The mesh size of a region will constrain the meshing of the neighbouring regions as conformal meshes are required. The coil mesh could be considered, but since it will be mostly isothermal, it should not require a fine mesh and we will not further constraint the mesh of other regions. The reported relative errors are with respect to the most refined fully resolved mesh.
@@
First, looking at the maximum temperature $T_{\mathrm{max}}$ on $\Omega_{\mathrm{TH}}$ as a function of time $t$ in Fig. 5, we can see a combination of the half-turns heating up due to the transient effects, and the cooling which is more pronounced at higher temperatures due to the Robin condition. For comparison, one simulation is shown (dotted) without cooling. As a result of competing effects, we have a maximum temperature of around $9$ K for this simulation at around $60$ s. Then, the system slowly cools down until the full current is reached (at $280$ s) and the transient effects fade away. The magnet then cools back down to $T_{\mathrm{He}}$. While doing so, the maximum temperature can appear at a different half-turn, which may be cooling down more slowly leading to ‘kinks’ in the curve, e.g. around $290$ s.
@@
Figure 6: (left) The relative error of the maximum temperature are calculated using a reference model, for the models in the legend and colour bar from Fig. 5. (right) Time average relative error of TSAs using different collar mesh sizes (colour, in meters) and TSA discretization sizes of the insulation (marker, in meters). The pole mesh size is half the collar mesh size.
@@
The relative error of $T_{\mathrm{max}}$, $\varepsilon_{\mathrm{rel}}T_{\mathrm{max}}$, with respect to a reference mesh, decreases with increasing number of DoFs, see Fig. 6. This is mainly related to the insulation mesh size, which is specified in Tab. 2. A peak in the relative error is observed at the end of the current ramp.
@@
We further noticed that the collar and pole mesh size does not influence the maximum temperature. The only noticeable difference is the local maximum temperature in the pole and collar region itself. This is not of great importance since aluminium alloy is very thermally conducting, and the different mesh sizes all result in low maximum temperatures with respect to the half-turn temperatures. Hence, this yields a similar cooling effect on the half-turns. This can be seen from the models with a coarser collar mesh which are still able to capture the maximum temperature accurately. However, when modelling a stainless steel collar, the collar exhibits larger temperature gradients and the collar mesh is more important to correctly describe the cooling.
@@ 3.1.2 TSA mesh
Next, we look at TSA models and the effect of the meshing parameters on the TSA solution. Unlike the fully resolved meshes, the TSA models employ quadrilateral elements to mesh the half-turns and wedges, as they are not constrained by a surrounding mesh and are approximately isothermal. Fig. 6 illustrates the mean error of a simulation for different collar and pole mesh sizes (assuming pole mesh size is half the collar mesh size) and different insulation mesh sizes, which determines the number of internal TSLs. The TSA is compared against the most refined fully meshed model.
By varying the meshing parameters for the TSA, we observed that the solution does not further converge toward the fully resolved reference mesh. This behaviour can partly be attributed to the modelling assumptions, as the two models are not necessarily expected to yield identical results. In addition, the discrepancy is likely related to the assumption of negligible tangential heat flux in the insulation layer. While a more sophisticated scaling factor might improve the accuracy, finding a general solution is likely cumbersome.
@@
Based on Fig. 6, it may seem that some TSA models yield errors much below $1\%$, however, more refined TSA models yield about $1.6\%$ mean error. Second, note that the insulation size only affects the number of DoFs, meaning few internal discretisations are sufficient to resolve the temperature gradient across the insulation layer. We conclude that the mean error depends on the collar and pole mesh sizes. For this reason, we select a single TSA model for the remainder of the discussion. This model is sufficiently converged in terms of mesh resolution while using the smallest number of DoFs, approximately 3 k.

### v2

@@ Abstract
Understanding the thermal behaviour of superconducting accelerator magnets is essential to ensure their stable and reliable operation. This work presents an extension of the Finite Element Quench Simulator (FiQuS) Multipole module to include collar and pole regions of accelerator magnets, which influences the overall thermal response. A thermal thin shell approximation (TSA), which is shown to be effective from previous works, is employed to model thermal insulation layers efficiently, replacing an insulation surface mesh. The main novelty of this work lies in the development of a method to model the thermal connection between the magnet winding and the collar and pole regions via the TSA.
To assess the accuracy and computational efficiency of this method, temperature and field variations are computed for a current ramp-up scenario. The thermal solution is coupled to a fully resolved magnetodynamic solution to capture the interaction between thermal and electromagnetic behaviour. The results obtained with the TSA are then compared to classical finite element (FE) solutions with explicitly meshed insulation domains. The TSA predicts the maximum temperature within 2–4 % of the reference solution while substantially reducing mesh complexity and achieving up to a 5 times speed-up in computation time. While the TSA has traditionally been employed for short-duration quench simulations with high heat fluxes between magnet turns, these results demonstrate its reliability and efficiency for current ramp scenarios with low heat fluxes, significantly expanding its application range beyond what has been previously reported in the literature.
@@ 1 Introduction
The thermal stability of superconducting magnets is of critical importance for particle accelerators using such technology. At the European Laboratory for Particle Physics (CERN), the Simulation of Transient Effects in Accelerator Magnets (STEAM) [1, 2] framework provides a collection of tools to model the transient behaviour of these magnets. Within this framework, the Finite Element Quench Simulator (FiQuS) [3] is a coupled magnetodynamic-thermal finite-element based tool used to model the behaviour of a variety of magnets types (e.g. high-temperature superconducting pancake coils [4], multipole magnets [3], canted cosine-theta magnets [5]). FiQuS relies on the Gmsh Python application programming interface [6] for geometry and mesh creation, and GetDP [7, 8, 9] for solving the governing equations.
Quench simulations typically involve large temperature gradients between adjacent half-turns, as the thermal conductivity of the superconducting cable is significantly larger than that of the surrounding electric insulation. Capturing these gradients accurately typically requires a high-quality, fully resolved mesh of the insulation layers [10, Sec. 5.3], leading to a large number of degrees of freedom (DoFs) and, consequently, high computational costs. For coupled FE simulations, two different meshes are often used for the thermal and magnetic models [11]. By enforcing suitable interface conditions, the thermal thin shell approximation (TSA) provides an efficient alternative without the need for explicit meshing [11, 12, 13, 14, 15].
Figure 1: Illustration of the electromagnetic computational domain $\Omega_{\mathrm{EM}}$ and thermal computational domain $\Omega_{\mathrm{TH}}$ for a 12 T dipole magnet design suggested in [16]. Only a quarter of the magnet is shown. The collar, pole, and ground insulation regions have been integrated into FiQuS. The materials are specified in Tab. 1.
While commercial tools like COMSOL Multiphysics can model thin thermal layers [17], these rely on proprietary, closed-source implementations. The detailed derivation of the TSA is instead presented and implemented in an open-source framework in [13]. Despite these advances, current state-of-the-art models still typically neglect the thermal influence of structural components such as the pole and collar.
Cooling mechanisms are critical for stable magnet operation. Standard approaches for modelling immersive cooling methods include convective cooling boundary conditions via a heat transfer coefficient [18], e.g. in FiQuS [19, Sec. 2.2], or by modelling a confined helium layer next to the insulation, as is done in ANSYS [20]. In the Large Hadron Collider (LHC) main dipoles, the non-impregnated Nb-Ti cables allow superfluid helium to permeate the coil and provide efficient cooling [21]. However, the amount of liquid helium required and its potential leakage can lead to a high cost.
@@ 2.1 Magneto-thermal formulation
Magnetodynamic equations are considered in the electromagnetic computational domain $\Omega_{\mathrm{EM}}$ in an $\mathbf{A}$-formulation. The magnetic vector potential $\mathbf{A}$ is related to the magnetic flux density $\mathbf{B}~=~\nabla\times\mathbf{A}$ and the electric field $\mathbf{E}~=~-\tfrac{\partial{\mathbf{A}}}{\partial{t}}~+~\nabla~\phi$. The electric scalar potential $\phi$ can be eliminated by a gauge choice: $\mathbf{A}^{*}~=~\mathbf{A}~+~\int_{t_{0}}^{t}~\nabla\phi\,~\mathrm{dt}$, leading to the strong $\mathbf{A}^{*}$-formulation with the governing differential equation [27] [28, Sec. 5.18]:
@@
where $\kappa=\kappa(|\mathbf{B}|,~T)$ denotes the thermal conductivity, $C_{V}=C_{V}(T)$ the volumetric heat capacity, $\vec{n}$ the normal to the boundary, and $Q_{\mathrm{ht}}$ the additional Joule losses in the half-turns, which are triggered when the local current density exceeds the current sharing threshold. This corresponds to Eq. (3) in [12]. The power density $P=\sigma|\mathbf{E}|^{2}$ is a modified, ramp rate dependent heat source, introduced to include a non-zero heat source below current sharing regime. Both the heat transfer coefficient $h=h(T)$ and the temperature of the liquid helium $T_{\mathrm{He}}=1.9$ K are used in Eq. (3) to define a Robin-type boundary condition for the cooling holes. The last equation imposes adiabatic conditions, i.e. no heat can escape through $\partial\Omega_{\mathrm{TH}}\!\setminus\!\Gamma_{\mathrm{Cool}}$. The half-turns region is modelled as a single homogeneous domain, where the different material properties are accounted for via volumetric fractions $f$, e.g. for $\sigma|_{\Omega_{\mathrm{ht}}}=f_{\mathrm{{}\mathrm{Cu}}}\sigma_{{}\mathrm{Cu}}$, as proposed in [12, Eqs. (1), (2) and (4)].
@@ 2.2.1 General idea and closest neighbour mapping
Figure 4: Illustration of a free closest neighbour map (left) and an enforced map (right). For clarity, only the collar temperature is shown at a specific point in time, for a quenching simulation and a stainless steel collar. Note the difference in local temperatures.
@@
For quench simulations or when using a low thermal conducting collar material, this closest neighbouring mesh node mapping might cause numerical artefacts in the simulation. This is illustrated by Fig. 4. On the left figure, two links are associated with the same mesh node, and since the total heat flux is conserved, there is an increase in local temperature compared to the figure on the right. This can lead to large discrepancies in maximum collar temperature. The problem can be avoided by applying a mortaring method to link both sides of a non-conformal mesh, enforcing equal heat flux by adding Lagrange multipliers [34]. This comes at the cost of additional unknowns and consequently a higher computational complexity.
Instead, a small mesh manipulation is implemented which enforces the mesh node position on the collar side to guarantee a one-to-one mapping. As illustrated by Fig. 4, this removes the artificial local differences in temperature, but compromises the collar mesh freedom.
@@ 2.2.2 Appropriate length scaling
As mentioned in the previous section, the TSA assumes that the tangential heat flux between different TSLs in the insulation layer is negligible. For simulations with fast quenching half-turns, this assumption leads to accurately computed maximum temperatures [12]. However, one must be cautious when generalising this. First of all, the average insulation thickness between the collar and half-turns or wedges (ground insulation in Fig. 1) is considerably larger than in-between half-turns and wedges (cable insulation in Fig. 1). As a result, more internal discretization levels in the TSA are needed. Moreover, the half-turn width ($w_{\mathrm{ht}}$ in Fig. 1) is typically much smaller than its height ($h_{\mathrm{ht}}$ in Fig. 4).
Taking into account the cable insulation with a thickness $d$, we make a distinction between lines along the radial and azimuthal direction. The former, appearing between one half-turn and the collar or pole, are assumed to have a contact length about $15\%$ larger ($h_{\mathrm{ht}}+2d]/h_{\mathrm{ht}}=1.1508$, see Fig. 4) for a 12 T dipole magnet. Similarly, the azimuthal lines, only appearing at the poles, represent a contact length which is $7\%$ larger ($w_{\mathrm{pole}}~/~w_{\mathrm{ht}}~=~1.0701$, see Fig. 1).
We use a scaling factor in order to diminish the error introduced by these geometric differences for the thin shell length, without compromising the computational speed by adding additional calculations or restraining the mesh. A similar scaling of material properties has previously been applied in [15] for pancake coils. In our case, we model the TSL as an extension of the half-turns (thus maintaining the bare half-turn length) and scale the material parameters with a factor, effectively modelling a longer thin shell, as depicted in Fig. 4. The main benefit of keeping the thin shell length the same as the half-turns on a geometrical level is for the mapping and for the mesh enforcement explained earlier. Concerning the TSLs connecting the wedges to the collar and pole regions, no correction is applied, assuming the insulation thickness is very thin compared to the wedge or half-turn length. This scaling law can only partially cover the effect of the tangential heat flow and it does not account for the mid-pole additional insulation. Moreover, since a linear scaling is applied to non-linear material properties, it cannot match the reference model perfectly. However, the approach leads to reasonable accuracy as will be shown in Sec. 3.
@@ 3.1.1 Fully resolved mesh models
The mesh quality is analysed by varying the mesh sizes, i.e. the characteristic size of the mesh triangles, and comparing the simulation results to those obtained with the most refined mesh. We consider variation of the insulation mesh and collar mesh sizes. The mesh size of a region will constrain the meshing of the neighbouring regions as conformal meshes are required. The coil mesh could be considered, but since it will be mostly isothermal, it should not require a fine mesh and further constrain the mesh of other regions. The reported relative errors are with respect to the most refined fully resolved mesh, shown in yellow in Fig. 5.
@@
First, looking at the maximum temperature $T_{\mathrm{max}}$ on $\Omega_{\mathrm{TH}}$ as a function of time $t$ in Fig. 5, we can see a combination of the half-turns heating up due to the transient effects, and the cooling which is more pronounced at higher temperatures due to the Robin condition. For comparison, one simulation is shown (dotted) without cooling, in which a thermal runaway occurs. In contrast, for the cooled simulations, these competing effects result in a maximum temperature of around $9$ K for this simulation at around $60$ s. Then, the system slowly cools down until the full current is reached (at $280$ s) and the transient effects fade away. The magnet then cools back down to $T_{\mathrm{He}}$. While doing so, the maximum temperature can move to a different half-turn, which may be cooling down more slowly leading to ‘kinks’ in the curve, e.g. around $290$ s.
@@
Figure 6: (left) The relative error of the maximum temperature compared to the reference model for various mesh sizes (colour bar of Fig. 5) and the TSA model with 2918 DoFs (black). (right) Time averaged relative error of TSAs using different collar mesh sizes (colour, in meters) and TSA discretization sizes of the insulation (marker, in meters). The pole mesh size is half the collar mesh size.
@@
The relative error of $T_{\mathrm{max}}$, $\varepsilon_{\mathrm{rel}}T_{\mathrm{max}}$, with respect to a reference mesh, decreases with increasing number of DoFs, see Fig. 6 (left). This is mainly related to the insulation mesh size, which is specified in Tab. 2. A peak in the relative error is observed at the end of the current ramp, after which artefacts in the relative error appear since the location of the maximum temperature changes over time. These artefacts are a consequence of the chosen error metric and can be safely disregarded.
@@
We further noticed that the collar and pole mesh size do not influence the maximum temperature. The only noticeable difference is the local maximum temperature in the pole and collar region itself. This is not of great importance since aluminium alloy is very thermally conducting, and the different mesh sizes all result in low maximum temperatures with respect to the half-turn temperatures. Hence, this yields a similar cooling effect on the half-turns. This can be seen from the models with a coarser collar mesh which are still able to capture the maximum temperature accurately. However, when modelling a stainless steel collar, the collar exhibits larger temperature gradients and the collar mesh is more important to correctly describe the cooling.
@@ 3.1.2 TSA mesh
Next, we look at TSA models and the effect of the meshing parameters on the TSA solution. Unlike the fully resolved meshes, the TSA models employ quadrilateral elements to mesh the half-turns and wedges, as they are not constrained by a surrounding mesh and are approximately isothermal. Fig. 6 (right) illustrates the mean error of a simulation for different collar and pole mesh sizes (assuming pole mesh size is half the collar mesh size) and different insulation mesh sizes, which determines the number of internal TSLs. The TSA is compared against the most refined fully meshed model.
By varying the meshing parameters for the TSA, we observed that the solution does not further converge toward the fully resolved reference mesh. This behaviour can partly be attributed to the modelling assumptions, as the two models are not necessarily expected to yield identical results. The discrepancy is likely related to the assumption of negligible tangential heat flux in the insulation layer. While a more sophisticated scaling factor might improve the accuracy, finding a general solution may be cumbersome.
@@
Based on Fig. 6, it can be seen that some TSA models yield errors much below $1\%$, however, the more refined TSA models yield about $1.6\%$ mean error. It can also be noted that the changing discretisation of the insulation does not alter the average error, meaning that few internal discretisations are sufficient to resolve the temperature gradient across the insulation layer. We conclude that the mean error depends mainly on the collar and pole mesh sizes. For this reason, we select a single TSA model for the remainder of the discussion. This model has sufficient mesh resolution while using the smallest number of DoFs, approximately 3 k.